# Risk Bounds for Mixture Density Estimation on Compact Domains via the $h$-Lifted Kullback–Leibler Divergence

### v1

@@ Appendix A The $h$-lifted KL divergence as a Bregman divergence
$D_{\phi}(p\,||\,q)$ $=\int_{X}\{\tilde{p}\log(\tilde{p})-\tilde{p}-1\}-\{\tilde{q}\log(\tilde{q})-\tilde{q}-1\}-\log(\tilde{q})(p-q)\mathrm{d}\mu$
$=\int_{\mathcal{X}}\tilde{p}\log(\tilde{p})-\tilde{q}\log(\tilde{q})-p\log(\tilde{q})+q\log(\tilde{q})\mathrm{d}\mu$
$=\int_{\mathcal{X}}\{p+h\}\log(\tilde{p})-\{q+h\}\log(\tilde{q})-p\log(\tilde{q})+q\log(\tilde{q})\mathrm{d}\mu$
@@
$=\int_{\mathcal{X}}\{p+h\}\log{\frac{p+h}{q+h}}\mathrm{d}\mu=\mathrm{KL}_{h}\left(p\,||\,q\right).$
Appendix B Experimental results
The code for all simulations and analyses of Experiment 1 and 2 is available in both the R and Python programming languages. The experiments and analyses presented were conducted in R. The figures were created in Python. The code repository is available here: https://github.com/hiendn/LiftedLikelihood.
Appendix C Technical results
@@
$\log N_{[]}(\mathscr{F},2\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq\log N(\mathscr{F},\varepsilon,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),$
for all $\varepsilon>0$.
@@
Let =$\mathscr{F}=\{f_{t}:t\in T\}$ be a function class satisfying
@@
Lemma 19 (Shalev-Shwartz and Ben-David (2014), Lem 26.7).
@@
Let $(X_{i})_{i\in[n]}$ be non-random elements of $\mathcal{X}$ and let $\mathscr{F}$ be a class of real-valued functions on $\mathcal{X}$. If $\varphi_{i}:\operatorname*{\mathbb{R}}\to\operatorname*{\mathbb{R}}$, $i\in[n]$, are functions vanishing at zero that satisfy for all $u,v\in\operatorname*{\mathbb{R}}$,
$\lvert\varphi_{i}(u)-\varphi_{i}(v)\rvert\leq\lvert u-v\rvert,$
then we have
@@
Lemma 22 (van der Vaart and Wellner, 1996, Lem. 2.3.1).

### v2

@@ C.1 The $h$-lifted KL divergence as a Bregman divergence
$D_{\phi}(f\,||\,g)$ $=\int_{X}\{\tilde{f}\log(\tilde{f})-\tilde{f}-1\}-\{\tilde{g}\log(\tilde{g})-\tilde{g}-1\}-\log(\tilde{g})(f-g)\mathrm{d}\mu$
$=\int_{\mathcal{X}}\tilde{f}\log(\tilde{f})-\tilde{g}\log(\tilde{g})-f\log(\tilde{g})+g\log(\tilde{g})\mathrm{d}\mu$
$=\int_{\mathcal{X}}\{f+h\}\log(\tilde{f})-\{g+h\}\log(\tilde{g})-f\log(\tilde{g})+g\log(\tilde{g})\mathrm{d}\mu$
$=\int_{\mathcal{X}}\{f+h\}\log{\frac{f+h}{g+h}}\mathrm{d}\mu=\mathrm{KL}_{h}\left(f\,||\,g\right).$
C.2 Proof of Proposition 2
Let $\tilde{f}=f+h$ and $\tilde{g}=g+h$. Since $h$ is positive, there exists some $\tilde{g}_{*}$ such that $\tilde{g}_{*}=\inf_{x\in\mathcal{X}}\left\{g(x)+h(x)\right\}>0$. Similarly, since $\mathcal{X}$ is compact, there exists some positive $\tilde{f}^{*}$ such that $0<\tilde{f}^{*}=\sup_{x\in\mathcal{X}}\left\{f(x)+h(x)\right\}<\infty$. Define $M=\sup_{x\in\mathcal{X}}\log\{\tilde{f}(x)/\tilde{g}(x)\}$. Then $M<\infty$, and
$\mathrm{KL}_{h}\left(f\,||\,g\right)=\int_{\mathcal{X}}\tilde{f}\log\frac{\tilde{f}}{\tilde{g}}\mathrm{d}\mu\leq\sup_{x\in\mathcal{X}}\log\frac{\tilde{f}}{\tilde{g}}\int_{\mathcal{X}}\tilde{f}\mathrm{d}\mu=2M<\infty.$
C.3 Proof of Proposition 3
Defining $\tilde{f}$ and $\tilde{g}$ as above, we have
The first inequality comes from the fundamental inequalities on logarithm $\log(x)\leq x-1$ for all $x\geq 0$. Indeed, let $f(x)=\log(x)-x+1$. We obtain $f^{\prime}(x)=\frac{1}{x}-1=\frac{1-x}{x}$. Then $f^{\prime}(x)<0$ if $x>1$ and $f^{\prime}(x)\geq 0$ if $x\leq 1$. Therefore, $f$ is strictly decreasing on $(1,\infty)$ and $f$ is strictly increasing on $(0,1]$. This leads to the desired inequality $f(x)\leq f(1)=0$ for all $x\geq 0$.
The next equality comes from the following identities:
$\int_{\mathcal{X}}\tilde{f}(\frac{\tilde{f}}{\tilde{g}}-1)\mathrm{d}\mu=\int_{\mathcal{X}}\frac{\tilde{f}^{2}-\tilde{f}\tilde{g}}{\tilde{g}}\mathrm{d}\mu=\int_{\mathcal{X}}\frac{\tilde{f}^{2}-\tilde{f}\tilde{g}-\tilde{f}\tilde{g}+\tilde{g}\tilde{g}}{\tilde{g}}\mathrm{d}\mu=\int_{\mathcal{X}}\frac{(\tilde{f}-\tilde{g})^{2}}{\tilde{g}}\mathrm{d}\mu=\int_{\mathcal{X}}\frac{(f-g)^{2}}{\tilde{g}}\mathrm{d}\mu.$
The last equality is followed from
$\int_{\mathcal{X}}\frac{-\tilde{f}\tilde{g}+\tilde{g}\tilde{g}}{\tilde{g}}\mathrm{d}\mu=-\int_{\mathcal{X}}\tilde{f}\mathrm{d}\mu+\int_{\mathcal{X}}\tilde{g}\mathrm{d}\mu=-\int_{\mathcal{X}}(f+h)\mathrm{d}\mu+\int_{\mathcal{X}}(g+h)\mathrm{d}\mu=-\int_{\mathcal{X}}h\mathrm{d}\mu+\int_{\mathcal{X}}h\mathrm{d}\mu=0.$
In fact, the proof of Proposition 3 follows the standard technique in the derivation of the estimation error, see for example Meir & Zeevi (1997).
Additionally, we can show that the $h$-lifted KL divergence satisfies a Pinsker-like inequality, in the sense that
$\sqrt{\mathrm{KL}_{h}\left(f\,||\,g\right)}\geq\mathrm{TV}(f,g),$
where $\mathrm{TV}$ represents the total variation distance between the densities $f$ and $g$. Indeed, this is easy to observe since
$\mathrm{KL}_{h}\left(f\,||\,g\right)$ $=\int_{\mathcal{X}}\left\{f+h\right\}\log\frac{f+h}{g+h}\mathrm{d}\mu=2\int\frac{f+h}{2}\log\frac{\left\{\frac{f+h}{2}\right\}}{\left\{\frac{g+h}{2}\right\}}\mathrm{d}\mu=2\mathrm{KL}\left(\frac{f+h}{2},\frac{g+h}{2}\right)$
$\geq 4\left\{\frac{1}{2}\int_{\mathcal{X}}\left|\frac{f+h}{2}-\frac{g+h}{2}\right|\mathrm{d}\mu\right\}^{2}\!\!\!=\left\{\int_{\mathcal{X}}\left|\frac{f+h}{2}-\frac{g+h}{2}\right|\mathrm{d}\mu\right\}^{2}\!\!\!=\left\{\frac{1}{2}\int_{\mathcal{X}}\left|f-g\right|\mathrm{d}\mu\right\}^{2}\!\!\!=\mathrm{TV}^{2}(f,g),$
where the inequality is due to Pinsker’s inequality:
$\sqrt{\frac{1}{2}\mathrm{KL}\left(f\,||\,g\right)}\geq\mathrm{TV}\left(f,g\right).$
C.4 Proof of Proposition 9
For choice (11), by the dominated convergence theorem, we observe that
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)$ $={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}$
$={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}.$
for every $p,q\in\mathcal{C}$ and $\pi\in\left(0,1\right)$, and thus
Similarly, for case (12), we have
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa_{n}((1-\pi)p+\pi q)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f(x_{i})+h(x_{i})}{\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)}$
$\qquad\qquad+\ \frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f\left(y_{i}\right)+h\left(y_{i}\right)}{\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)}$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(x_{i}\right)-q\left(x_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)\right]^{2}}+\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(y_{i}\right)-q\left(y_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)\right]^{2}}.$
@@ C.4 Proof of Proposition 9
$\sup_{p,q\in\mathcal{C},\pi\in\left(0,1\right)}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq\frac{2c^{2}}{a^{2}}<\infty,\text{ as required.}$.
Appendix D Technical results
@@
$\log N_{[]}(\mathscr{F},2\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq\log N(\mathscr{F},\varepsilon,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),\text{ for all $\varepsilon>0$.}$
@@
For some metric $d$ on $T$, let =$\mathscr{F}=\{f_{t}:t\in T\}$ be a function class:
@@
Lemma 21 (Shalev-Shwartz & Ben-David (2014), Lem 26.7).
@@
Let $(X_{i})_{i\in[n]}$ be non-random elements of $\mathcal{X}$ and let $\mathscr{F}$ be a class of real-valued functions on $\mathcal{X}$. If $\varphi_{i}:\operatorname*{\mathbb{R}}\to\operatorname*{\mathbb{R}}$, $i\in[n]$, are functions vanishing at zero that satisfy for all $u,v\in\operatorname*{\mathbb{R}}$, $\lvert\varphi_{i}(u)-\varphi_{i}(v)\rvert\leq\lvert u-v\rvert,$ then we have
@@
Lemma 24 (van der Vaart & Wellner, 1996, Lem. 2.3.1).